# LLMs Homogenize Values in Constructive Arguments on Value-Laden Topics

### v1

@@ 4.3. RQ3: What Types of Comments Do People Find More Aligned with Their Values?
We collected 1,122 evaluations from a total of 180 participants: who reviewed randomly selected pairs of human-written and LLM-rewritten comments. We found that on average both Indian (59.9%) and American participants (61.1%) reported LLM-rewritten comments to be more aligned with their values than human-written ones. We conducted multiple Chi-square tests with Bonferroni correction to examine if the types of comment participants found aligned with their values differed based on their stances on same-sex marriage and Islam. Table 6 shows the distributions of participants’ stances across these issues.
@@
American Participants. We found that the type of comment Americans found more aligned with their values differed significantly based on their stances on homophobia ($\chi^{2}(2,N=296)=20.51,p<0.0001,\phi=0.26$); but with a small effect size (see Figure 7). Post hoc analyses with Bonferroni corrections showed that Americans who opposed same-sex marriage (56.9%) found human-written comments significantly more aligned with their values than corresponding LLM-rewritten versions. They shared that their chosen comments better reflected their values around traditional ways of marriage between man and woman. An American participant wrote:
@@
Similarly, we observed a significant difference in American participants’ preferences for different comments based on their stances on Islam (see Figure 7), with a small effect only ($\chi^{2}(2,N=272)=13.92,p<0.001,\phi=0.23$). Americans who opposed Islam selected human-written comments significantly more than expected (53%) for being aligned with their values. They commented that their chosen comments accurately captured their perceptions of Islam by highlighting issues like terrorism, poor treatment of women and LGBTQ communities, lack of reform, and misalignment with Western liberal values. Some participants also pointed that the comments they did not choose were more woke, rambling, less factual, less realistic, and read like generic AI. An American participant shared:
@@
Potential Reasons Behind Participants’ Preferences for Different Comments. In 4.2 (see Table 4), we observed that LLM downplayed Traditional values and emphasized equality and tolerance (Universalism) towards same-sex marriage when rewriting comments from participants, who were against such marriage. This shift in value might explain why Indian and American participants, opposing same-sex marriage found human-written comments more representative of their traditional viewpoints on marriage.
Similarly, LLM downplayed Security concern around extremism and highlighted tolerance (Universalism) of diverse religions when rewriting constructive comments on Islamophobic thread (see Table 4). These changes in value framing might have impacted American participants to choose human-written comments that captured their criticism of Islam well. Moreover, as noted in 4.2, for comments opposing same-sex marriage and Islam, the value shifts introduced by LLM often changed their stance, making the rewritten versions more neutral or supportive. This may also explain why participants with opposing views felt human-written comments were more aligned with their values.
On the other hand, LLM did not significantly change underlying value expressions when rewriting comments which were supportive of same-sex marriage and Islam (see Table 4). This might explain why Indian participants with favorable stance on these cases, did not find one type of comment more aligned with their values than the other. However, the American participants supporting same-sex marriage and Islam preferred LLM-rewritten comments significantly more than the actual human-written comments. This might be because more American participants held supportive views of same-sex marriage and Islam than the Indians (see Table 6).
@@ 5.1. Conservative Bias and Prosocial Leanings of LLMs
Yet, framing value homogenization simply as bias or underperformance risks obscuring the deeper issue (Neuman et al., 2025; Hagendorff, 2025). Alignment is never a purely technical process but an exercise in normative judgment about what counts as harm, fairness, or bias. Current alignment practices tend to privilege ethical frameworks more closely associated with progressive traditions (Glaese et al., 2022; Bai et al., 2022), thereby reinforcing a particular vision of “prosociality.” The dilemma, then, is whether alignment should prioritize ethical standards that safeguard vulnerable communities (e.g., racial, gender, religious, or ethnic minorities) or whether it should accommodate a broader spectrum of socio-cultural, political, and moral values, even when those may conflict with the liberal notion of harm-reduction goals.
@@ 5.2. Impact on Online Discourse on Value-Laden Topics
These asymmetries can contribute to “quality-of-service harms” (Shelby et al., 2023), as models may systematically underserve users with Conservative value systems, potentially deepening mistrust, exclusion, and alienation in digital spaces. In fact, research shows that undesirable exposure to ideologies that do not align with one’s values (Bai et al., 2022) and experiences of socio-political alienation can drive conservatives toward more polarized positions (Aberbach, 1969; McClosky and Chong, 1985; Abcarian and Stanage, 1965) and far-right platforms such as Truth Social (Brown and Baggili, 2023). Thus, LLM-driven prosocial value homogenization might lead to uneven benefits and inadvertently push users to more radical viewpoints.
@@ 5.3. Limitations and Future Work
Our work has some limitations. First, we focused on two value-laden topics: Islamophobia and homophobia. Thus, our findings may not generalize to other value-laden topics, such as abortion, gun regulation, climate change, or vaccination. Future work should explore if value homogenization by LLMs extends to diverse value-laden topics. Second, in our study we only experimented with GPT-4. Although prosocial leanings have been documented across multiple models (Neuman et al., 2025; Shen et al., 2024; Russo et al., 2025), further work is needed to examine whether different models also shift values when performing constructive rewriting. Third, our study recruited participants only from India and the United States. The absence of observed cross-cultural differences in constructive comments on value-laden topics may be due to the limited sample size. Future work should examine how value homogenization unfolds across a wider range of cultural contexts and broad samples. Finally, we only focused on constructive writing and value expressions in English. Research shows that while LLMs tend to produce similar value orientations across many European languages (e.g., French, German), outputs can diverge sharply when comparing languages from more distant cultural and linguistic traditions, such as English and Chinese (Cahyawijaya et al., 2024). Future work should investigate value homogenization across multilingual contexts.
@@ 6. Conclusion
Through experiments with 347 participants from India and the US, this work unpacks how LLMs contribute to value homogenization when facilitating constructive discourse on sensitive, value-laden topics. Our findings show that LLMs systematically downplay Conservative values, foreground prosocial values, and at times change the stance in comments from opposing to supporting on value-laden topics. These dynamics raise pressing ethical tensions and moral dilemmas for the deployment of LLMs in cross-cultural, value-sensitive contexts. Addressing these challenges would require our socio-technical systems to be more attuned to and mindful towards diverse human perspectives on contested value-laden issues.
@@ 2. Islamophobia is a great thing
Feature Statistics Human-written comment (Average) LLM-rewritten comment (Average)
Length W=3009.5, Z=-9.1, p<0.000001, r=0.44 62.1 80.1
Readability W=167.5, Z=-12.37, p<0.000001, r=0.6 7.4 15.3
Named entity W=4621.5, Z=3.95, p<0.00001, r=0.19 1.9 1.5
Discourse connective W=6811, Z=-2.86, p<0.01, r=0.14 5.99 6.61
Stance adverbial - 0.59 0.40
Reasoning verbs and modals W=1355, Z=-10.48, p<0.000001, r=0.51 4.5 7.34
Root clauses W=2423, Z=-8.34, p<0.000001, r=0.41 3.5 4.6
Politeness W=6161.5, Z=-4.74, p<0.00001, r=0.23 12.73 14.73

### v2

@@ 4.3. RQ3: What Types of Comments Do People Find More Aligned with Their Values?
We collected 1,122 evaluations from a total of 180 participants: who reviewed randomly selected pairs of human-written and LLM-rewritten comments. We found that on average both Indian (59.9%) and American participants (61.1%) reported LLM-rewritten comments to be more aligned with their values than the human-written ones. We conducted multiple CHI-square tests with Bonferroni correction to examine if the types of comment participants found aligned with their values differed based on their stances on same-sex marriage and Islam. Table 9 shows the distributions of participants’ stances across these issues.
@@
American Participants. We found that the type of comment Americans found more aligned with their values differed significantly based on their stances on homophobia ($\chi^{2}(2,N=296)=20.51,p<0.0001,\phi=0.26$); but with a small effect size (see Figure 8). Post hoc analyses with Bonferroni corrections showed that Americans opposing same-sex marriage (56.9%) found human-written comments significantly more aligned with their values than corresponding LLM-rewritten versions. They shared that their chosen comments better reflected their values around traditional ways of marriage between man and woman. An American participant wrote:
@@
Similarly, significant difference exists in American participants’ preferences for comments based on their stances on Islam (see Figure 8), with a small effect only ($\chi^{2}(2,N=272)=13.92,p<0.001,\phi=0.23$). Americans who opposed Islam selected human-written comments significantly more than expected (53%) for being aligned with their values. They commented that their chosen comments accurately captured their perceptions of Islam by highlighting issues like terrorism, poor treatment of women and LGBTQ communities, lack of reform, and misalignment with Western liberal values. Some participants also pointed that the comments they did not choose were more woke, rambling, less factual, less realistic, and read like generic AI. An American participant shared:
@@
Potential Reasons Behind Participants’ Preferences for Different Comments. In 4.2 (see Table 4.2), we observed that LLM downplayed Conservative values (Tradition, Conformity, Security concern) and emphasized equality and tolerance (Universalism and Benevolence) towards same-sex marriage when rewriting comments from participants, who were against such marriage. This shift in value might explain why Indian and American participants, opposing same-sex marriage found human-written comments more representative of their traditional viewpoints on marriage.
Similarly, LLM downplayed Security concern around extremism and highlighted tolerance (Universalism) of diverse religions when rewriting comments on Islamophobic thread (see Table 4.2). These changes in value framing might have influenced American participants to choose human-written comments that captured their criticism of Islam well. Moreover, as noted in 4.2, for comments opposing same-sex marriage and Islam, LLM-mediated value shifts often changed their stance, making the rewritten versions more neutral or supportive. This may also explain why participants with opposing views felt human-written comments were more aligned with their values.
In contrast, when rewriting comments supportive of same-sex marriage and Islam, the LLM either preserved the original values or further amplified Benevolence (concern for others) (see Table 4.2). This might explain why Indian participants who held supportive views on these issues did not show preference for either type of comment. However, American participants who supported same-sex marriage and Islam preferred the LLM-rewritten comments significantly more than the original human-written ones. This difference may stem from a larger proportion of American participants holding supportive views on these issues compared to Indian participants (see Table 9).
@@ 5.1. Conservative Bias and Prosocial Leanings of LLMs
Yet, framing value homogenization simply as bias or underperformance risks obscuring the deeper issue (Neuman et al., 2025; Hagendorff, 2025). Alignment is never a purely technical process but an exercise in normative judgment about what counts as harm, fairness, or bias. Current alignment practices tend to privilege ethical frameworks more closely associated with progressive traditions (Glaese et al., 2022; Bai et al., 2022), thereby reinforcing a particular vision of “prosociality.” However, there are notable exceptions. For instance, political ideologies have contributed to deliberate censorship of model outputs, as in the case of DeepSeek (Wang, 2025; Booth and DanMilmo, 2025), or pushed systems like Grok toward more right-leaning and antisemitic outputs (Wirtschafter and Nadgir, 2025; Folco, 2025). The dilemma, then, is whether alignment should prioritize ethical standards that safeguard vulnerable communities (e.g., racial, gender, religious, or ethnic minorities) or whether it should accommodate a broader spectrum of socio-cultural, political, and moral values, even when those may conflict with the liberal notion of harm-reduction goals.
@@ 5.2. Impact on Online Discourse on Value-Laden Topics
These asymmetries can contribute to quality of service harms (Shelby et al., 2023), as models may systematically underserve users with Conservative value systems, potentially deepening mistrust, exclusion, and alienation in digital spaces. In fact, research shows that undesirable exposure to ideologies that do not align with one’s values (Bai et al., 2022) and experiences of socio-political alienation can drive conservatives toward more polarized positions (Aberbach, 1969; McClosky and Chong, 1985; Abcarian and Stanage, 1965) and far-right platforms such as Truth Social (Brown and Baggili, 2023). Thus, LLM-driven prosocial value homogenization might lead to uneven benefits and inadvertently push users to more radical viewpoints.
@@ 5.3. Limitations and Future Work
Our work has some limitations. As an exploratory first step, we compared participants from two different cultures: India (collectivisitc culture) and the United States (individualistic culture). Thus, our findings on what values people highlight when writing constructively on value-laden topics may not generalize to diverse cultural contexts, including subcultures in these countries. Future research should explore how these dynamics manifest across different populations. Second, our samples are not nationally representative because we used Prolific’s Harmful Content Prescreener, which limited recruitment to participants who had self-identified as being comfortable with reviewing sensitive material. Moreover, Prolific does not offer representative sampling for India, which further constrained our options.
Third, we focused on two value-laden topics: Islamophobia and homophobia that are salient to socio-political and moral rhetoric both in India and the US. While we acknowledge that our findings may not generalize to other value-laden topics, such as abortion, gun regulation, climate change, or vaccination, capturing the full spectrum of human values across diverse issues and cultures is beyond the scope of this study. Rather we view this work as an initial step toward understanding how LLMs engage with and homogenize value expressions across cultural contexts. We hope future research will build on this foundation to examine whether similar patterns emerge on other value-laden issues.
Fourth, in our study we only experimented with GPT-4. Although prosocial leanings have been documented across multiple models (Neuman et al., 2025; Shen et al., 2024; Russo et al., 2025), further work is needed to examine whether different models also shift values when performing constructive rewriting. Finally, we only focused on comment writing and value expressions in English. Research shows that while LLMs tend to produce similar value orientations across many European languages (e.g., French, German), outputs can diverge sharply when comparing languages from more distant cultural and linguistic traditions, such as English and Chinese (Cahyawijaya et al., 2024). Future work should investigate value homogenization across multilingual contexts.
@@ 6. Conclusion
Through experiments with 465 participants from India and the US, this work unpacks how LLMs contribute to value homogenization when facilitating constructive discourse on sensitive, value-laden topics. Our findings show that LLMs systematically downplay Conservative values, foreground prosocial values, and at times change the stance in comments from opposing to supporting on value-laden topics. These dynamics raise pressing ethical tensions and moral dilemmas for the deployment of LLMs in cross-cultural, value-sensitive contexts. Addressing these challenges would require our socio-technical systems to be more attuned to and mindful towards diverse human perspectives on contested value-laden issues.
Acknowledgements.
We thank Cornell Global AI Initiative and Infosys for supporting this work.
@@ 2. Islamophobia is a great thing
Table 11. Example codes with participant’s open-ended responses along with the higher-level themes and distributions in the dataset.
Theme Code Description Example
Personal
value
84%
responses
Universalism
selected
comment
based on
values
The comments I selected were well-reasoned
and align with my beliefs because I support
gay marriage. I believe in an individuals right
to choose how to live their life.
Linguistic
property
12%
responses
Short, direct
argument
selected
comment
based on
language
I went with the ones that were shorter, more
direct and to the point. I think those will be
understood better than the longer ones.
Extreme
opposition
4%
responses
Non-
alignment
no comment
aligned
with
participant’s
values
Actually, not a single one aligned
themselves with my values. I believe Islam
is evil! Woman forced to cover themselves
from head to toe or brutally murdered by
their families! I do not give that so called
religion any mercy what so ever.